OpenStudy (anonymous):

Sarah draws a card from a deck of 52 cards. She receives 40 cents for a heart, 50 cents for an ace, and 90 cents for the ace of hearts. The cost of a draw is 15 cents. How much will Sarah lose if she plays this game?

OpenStudy (amistre64):

if I played Id lose 15 cents

OpenStudy (anonymous):

thats what I thought but when i put that as my anwser it said it was wrong

OpenStudy (amistre64):

how many draws she get?

OpenStudy (anonymous):

it doesnt say thats all the info i got

OpenStudy (anonymous):

did you try 0?

OpenStudy (amistre64):

well then, the most she could lose is 15 cents..... but maybe those aint vegas odds :)

OpenStudy (anonymous):

no I havent I have one more try to get it right you get four and i used them all up

OpenStudy (amistre64):

your allowed to miss a few questions right? :)

OpenStudy (anonymous):

haha yea i missed one already lol i guess i could miss this one i have two more questions that im about to post if you could help me with them that would be great

OpenStudy (amistre64):

i aint none to good with probability yet..... but i could take a gander at 'em

OpenStudy (anonymous):

there is something wrong with this question this is the solution i got...E = 40(12/52)+50(3/52)+90(1/52) = 13.8. This means Sarah expects to win 13.8 cents. However, she pays 15 cents to play the game. Therefore, she actually loses money: 13.8-15 = 1.2 cents.

OpenStudy (amistre64):

why would sarah expect to win a game? i agree, that question is bad...maybe even evil..... lol

OpenStudy (anonymous):

can it be 40(1/4)+50(1/13)+90(1/52) ?

OpenStudy (anonymous):

okay I understand it now just talked to someone in class about it...because there's the option for the ace of hearts you have to take that out of the other options,....you cant win both cash prices at once